Ruth Lynn Seltzer

80 Orchard St. Medford, MA 02155 (617) 395–8244

Education

Wellesley College Wellesley, MA Sept. 1986– June 1990

Bachelor of Arts Degree in Psychology. Minor in Computer Science.

Computer Skills

Operating Systems 5 years MS-DOS, 4 years Unixtm (BSD 4.3, ULTRIX 4.0), 1 year using MIT Project Athena, 3 years VMS, 2 years TOPS-20, 3 years CP/M, 5 years Apple-DOS.

Hardware Platforms IBM PC (and compatibles), DEC VAX, IBM RT, DECSystem 20, Apple II, DEC Rainbow, DEC PDP-11, Amiga.

Languages Pascal, BASIC, 6502 Assembly, VAX Assembly, LOGO.

Software Packages Lotus 123, SuperCalc, VisiCalc, Dbase, Wordperfect, DisplayWrite, WordStar, Scribe, Latex, emacs, edt, DigiPaint, Microsoft Windows, SPSSX.

Communications TCP/IP, DecServer, Internet mail, Kermit.

Work Experience

March  Badger Engineers, Inc.  Cambridge, MA

1991-Present Project Secretary for McNeil Technical Center Project.

Secretary for IFG Purchasing Department.

1990-Feb 1991 ECCO-Staffbuilders Temporary Agency Cambridge, MA Assignments completed:

Badger Engineering – Word Processing Department. Badger Engineering – Project Secretary for Chevron PAR Project.

1987-1990 Massachusetts Institute of Technology Cambridge, MA

Wrote manuals for staff training and introductory guides for faculty and student use. Supported computers used for Electrical Engineering and Computer Science Department courses and faculty.

Ruth Lynn Seltzer

Page 2

Summer 1989 Creative Temporary Agency Plainview, NY

Assignments completed:

Fairchild Weston – Compiled documention for computerized weapon

training systems.

NEC America – General office support. Lincoln Savings Bank – Mortgage Processing. Loral – Accounting.

Jerome I. Silverman Company, CPA.

{Summer 1989, \atop 1987} Plainview, NY

Developed and documented database software to track incoming mail, sales and inventory, and trained personnel on use of these programs. Prepared a computerized forecasting model for profits and losses and documented how model could be extended.

Summer 1986 Syosset Central School District Syosset, NY

Managed stage production of summer theatrical production. Instructed students in set construction.

Participated in summer theatrical productions as stage crew member.

{Summer 1983– \atop 1985}

Skills

Able to read Spanish.

Activities

Designed posters for M.I.T. Lecture Series Committee. Competed at the 1989 National Unicycle Meet. Tutored for Computer Science at Wellesley College.